OpenStudy (anonymous):

If my age is divided by 5, the remainder is 4. If it’s divided by 3, the remainder is 2. If it’s divided by 7, the remainder is 5. How old am I?

OpenStudy (anonymous):

x/5 =4 x/3 =2 x/7=5 x/5 + x/3 + x/7 =4+2+5 x(1/5 + 1/3 + 1/7)=11 x=11(1/5 + 1/3 + 1/7)

OpenStudy (anonymous):

what is the value of x?

OpenStudy (anonymous):

you can do what remains

OpenStudy (anonymous):

on no. try chinese remainder theorem

OpenStudy (anonymous):

will work it out if you like

OpenStudy (anonymous):

I agree with satellite; the working above in nonsense.

OpenStudy (anonymous):

is*

OpenStudy (dumbcow):

i believe its 89

OpenStudy (anonymous):

That is meaningless; like above, use the chinese remainder theorem.

OpenStudy (anonymous):

i believe that it is early in the morning to use crt, but it is actually interesting to list. remainder of 4 when divided by 5 9, 14, 19, 24, 29, ... and so see what happens when you consider the remainders when dividing by 3. the patter in 0, 2, 1, 0, 2 , 1,

OpenStudy (anonymous):

in any case method is described clearly here http://marauder.millersville.edu/~bikenaga/numbertheory/chinese-remainder/chinese-remainder.html

OpenStudy (anonymous):

for a really nice description of the history of the chinese remainder theorem see "the mathematical experience"

OpenStudy (anonymous):

so is the answer 89 ?

OpenStudy (anonymous):

\[x = 89 + 105k \] But realistically, k =0 is the only solution.

OpenStudy (anonymous):

@INewton : what is k ?

OpenStudy (anonymous):

k is an integer where \[k \geq 0 \]

OpenStudy (anonymous):

\[x \equiv \begin{cases} 2\mod\ 3 \\ 4\mod\ 5 \\ 5\mod\ 7\end{cases}\] \[\text{Let } x_1 \equiv \begin{cases} 1\mod\ 3 \\ 0\mod\ 5 \\ 0\mod\ 7\end{cases} \implies x_1 \equiv \begin{cases} 1\mod\ 3 \\ 0\mod\ 35 \end{cases} \implies x_1=35x_1' \text{ for some integer } x_1'\] \[\text{Therefore } 35x_1' \equiv 1 \mod 3 \implies 2x_1' \equiv 1\mod 3 \] \[\text{Let's take } x_1' = 2 \implies x_1 = 70\] Following the obvious notation we find similarly: \[x_2 = 21 \text{ and } x_3 = 15 \] \[\text{Clearly, } 2x_1 + 4x_1 + 5x_3 = x_{min} + 105k \implies x_{min} = 299 - 105k\] For the smallest positive integer x we have x = 89 (k=2), be we also have solutions of the form \[x = 89 + 105k\ \forall k \in \mathbb{N} \]

OpenStudy (anonymous):

Ignore the typos ¬_¬ (e.g. it should be 2x_1 + 4x_2 + 5x_3)